OpenStudy (anonymous):

If f of x equals x + 2 over x + 8 and 2x + 1 over x + 8, then find f plus g of x

OpenStudy (anonymous):

\[f \left( x \right) = \frac{ x+2 }{ x+8 } and \frac{ 2x+1 }{ x+8 }, then find f+g \left( x \right)?\]

OpenStudy (anonymous):

yeah a. 5x=4 b. x-4 c. x=14 d. 5x-4